A farmer had a herd of 24 animals, some sheep and some goats. He sold the sheep
for $140 each and the rest of the herd, the goats for $100 each. If he had sold the whole herd at the uniform price of $120 per animal, he would have gained 360
more dollars. How many goats did he sell?

Answer by vksarvepalli(154)

let the no. of sheep be s
and no. of goats be g
so we have s+g=24 -------------- 1
also if he had sold the whole herd at the uniform price of $120 per animal
he would have earned 24*120=2880$
this 2880$ is 360 more than what he actually earned
so he earned 2880-360 = 2520$
i.e., 140*s + 100*g = 2520 ----------2
from equations 1 & 2
40s = 2520-2400 = 120
so s = 3
and g = 21
thus the no. of goats he sold = 21 Ans